OpenStudy (anonymous):

Some how I got knocked off the system: emccormick the question is the radical expression of 17^1/6

OpenStudy (anonymous):

There you are somehow I was bumped off the system, glad to see you back!

OpenStudy (e.mccormick):

To review: The 4th root jumps up to an exponent as \(\frac{1}{4}\) \[\Huge \sqrt[4]{U}=U^{\frac{1}{4}}\] And inversely: \[\Huge U^{\frac{1}{4}}=\sqrt[4]{U}\] \(\frac{1}{4}\) jumps down as 4 when becoming a root. So, what happes to your \(\Huge 17^{\frac{1}{6}}\)?

OpenStudy (anonymous):

Thank you, this time I copied and pasted in Word for my notes to read over later!

OpenStudy (anonymous):

I thought it was missing th

OpenStudy (anonymous):

the symbols.

OpenStudy (anonymous):

How is this written as a radical expression?

OpenStudy (e.mccormick):

I am using \(\LaTeX\) to do all this.

OpenStudy (anonymous):

What is that?

OpenStudy (e.mccormick):

The thing that lets me format it on this page.

OpenStudy (anonymous):

This is only my 3rd time using this format so I don't know much about it yet but when I can I will do a review.

OpenStudy (anonymous):

I would have just used the radical symbol and written the ^1/6(17).

OpenStudy (e.mccormick):

Not quite. Something happens to the 1/6th.

OpenStudy (anonymous):

Oh yeah you said it changes to 1/4

OpenStudy (e.mccormick):

No, 4 and 1/4 are related. 1/6 becomes something related to 1/6th.

OpenStudy (anonymous):

ok,

OpenStudy (e.mccormick):

Here are a couple references for you that show more of these: http://www.regentsprep.org/Regents/math/algtrig/ATO1/FractionalExp.htm http://www.purplemath.com/modules/exponent5.htm That regentsprep one is really good on this topic.

OpenStudy (anonymous):

Great, I copied and pasted it into my notes. What do we do with the 1/6?

OpenStudy (e.mccormick):

Invert it.

OpenStudy (anonymous):

So is ti written as a square root with the symbol as ^6/1(17)

OpenStudy (e.mccormick):

Basically.... but what is ANYTHING when it is put over 1? Say: \(\cfrac{x}{1}\) That 1 does not matter, right? \(\cfrac{x}{1}=x\)

OpenStudy (anonymous):

So the it would be ^6√17

OpenStudy (e.mccormick):

Yes.

OpenStudy (anonymous):

Thank you, I am actually learning more here than in the classroom!

OpenStudy (e.mccormick):

And that is the relationship between roots and exponents in a nutshell. Sometimes you don't have time to ask the questions in class. Here, you can ask away.

OpenStudy (anonymous):

You are right! Thank you for being here. I am going to figure out what do do for dinner now! Hope to see you here again!

OpenStudy (e.mccormick):

Have fun!

OpenStudy (e.mccormick):

FYI: When you want to try and get the attention of someone, at tag them. For example: @Tuiti See how that sent you a message? Sometimes people are offline or busy, but in cases like that post that you got bumped and could not find again, the tag would have told me where you were.

OpenStudy (anonymous):

Great tip. I did search first but could not find your name! LOL! THANKS!